Titanium is used in airplane bodies because it is strong and light. It has a density of 4.55 g/cm^(3). If a cylinder of titanium is 7.75 cm long and has a mass of 153.2 g, calculate the diameter of the cylinder.


1
Expert's answer
2020-09-15T13:54:38-0400

The formula for volume of a cylinder is V = "\pi"r2h.

The volume can be found from the given mass and density:

V = m / "\rho" .

Therefore,

r2 = V / "\pi"h = m / "\pi"h"\rho" = 153.2g / (3.14 * 7.75cm * 4.55g/cm3) = 1.3836

r = 1.1763 cm

d = 2r = 2.35 cm

Answer: 2.35 cm